# Expanding the reach of diffusing wave spectroscopy and tracer bead microrheology

### v1

@@ Abstract
Diffusing Wave Spectroscopy (DWS) is an extension of standard dynamic light scattering (DLS), applied to soft materials that are turbid or opaque. The propagation of light is modeled using light diffusion, characterized by a light diffusion coefficient that depends on the transport mean free path $\ell^{\ast}$ of the medium. DWS is highly sensitive to small particle displacements or other local fluctuations in the scattering properties and can probe subnanometer displacements. Analyzing the motion of beads in a viscoelastic matrix, known as one-bead microrheology, is one of the most common applications of DWS. Despite significant advancements since its invention in 1987, including two-cell and multi-speckle DWS, challenges such as merging single and multi-speckle data and limited accuracy for short correlation times persist. Here, we address these issues by improving the two-cell Echo DWS scheme. We propose a calibration-free method to blend and merge Echo and two-cell DWS data and demonstrate the use of regularized inversion algorithms to enhance data quality at very short times. Building on this, we introduce stable corrections for bead and fluid inertia, significantly improving the quality of microrheology data at high frequencies.
@@ I Introduction
Diffusing Wave Spectroscopy (DWS) is a standard light scattering technique used to study the dynamics of colloids, colloid-polymer mixtures, surfactant solutions and other soft matter systems that display interesting complex relaxation dynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Introduced in 1987 [14, 15] DWS is an extension of the standard dynamic light scattering (DLS) approach [16]. In contrast to DLS it is applied to materials that are turbid or opaque. In such materials, light is not scattered only once, as is the case in DLS, but multiple times. The propagation of light is often modeled using light diffusion characterized by a light diffusion coefficient that depends on the transport mean free path $\ell^{\ast}$ of the medium. Coherent light becomes dephased due to the motion of the scatterers in the medium [17, 18, 19], which can be quantified by analyzing the Intensity time auto-Correlation Function (ICF) of scattered light, analogous to the case of DLS.
@@
DWS is commonly applied to samples such as dense suspensions and emulsions [20, 21, 22, 12, 23, 11, 24]. Alternatively, it can be applied to transparent samples by adding tracer beads that sense the dynamic environment of the medium [25, 26, 7, 5, 27, 28, 29, 22, 13]. The advantage of DWS over DLS is that it can be applied to turbid media that are not accessible to classical DLS due to dominant multiple light scattering [30]. Additionally, DWS is highly sensitive to small particle displacements or other local fluctuations of the scattering properties. DWS can probe subnanometer displacements, as demonstrated by Weitz et. al. in 1989 shortly after the method was introduced [31]. The high sensitivity of DWS is especially useful when tracer beads are embedded in a viscoelastic medium of interest. DWS can measure viscoelastic solids with elastic moduli up to 100 kPa or more [32] and can study relaxations on microsecond timescales, such as Rouse-Zimm and bending modes of polymer dynamics [33, 6, 7].
@@
Analyzing the motion of beads in a viscoelastic matrix is known as (single) bead microrheology [5, 26, 28], which is one of the most common applications of DWS. Pioneering work by Mason, Weitz, and others showed that the mean squared displacements of the beads $<\Delta\vec{r}^{2}(t)>$ can be directly converted into the linear viscoelastic spectrum, or complex modulus $G^{\ast}(\omega)=G^{\prime}(\omega)+iG^{\prime\prime}(\omega)$, of the materials if the bead radius is known [25, 34, 35]. DWS provides access to this information from the sub-microsecond scale to seconds and minutes, opening up opportunities in rheology to study viscoelastic moduli over an enormous rage of frequencies with a single measurement [7]. Several reviews of the theory underlying the method, along with numerous examples of its applications, have been published. We refer the reader to the literature for further details [1, 3, 4, 5, 27, 28, 13].
@@
Since the invention of DWS in 1987, several substantial improvements to the method have been made. The original approach relied on taking time averages of the intensity correlation function (ICF), which requires the sample to be in a liquid state. Two-cell DWS, introduced in 2000 [36], and other ensemble averaging approaches [37, 38] provided access to short correlation times for solid, nonergodic systems. Multispeckle DWS, using either a camera or the Echo-scheme, was introduced in the early 2000s and it provides access to longer correlation times for highly viscous or viscoealstic solids such as glasses and gels [39, 40, 41, 42, 43, 44]. In 2017, an automated method for determining the optical transport and absorption mean free paths for absorbing and non-absorbing samples was developed [29]. These advancements have made it possible to access the ICF over a range from typically 10 ns to tens of seconds or more while keeping the measurement time on the order of a few minutes. Only when studying even longer relaxation times does the measurement time need to be further increased. Despite these significant advances, shortcomings in the implementation of these methods can still lead to errors or reduced data quality. Moreover, access to very short correlation times is hindered by photon shot noise and the need to normalize the ICF, which, in its current implementation, introduces bias.
@@
In the present work, we address these problems and propose two substantial improvements to the implementation of the two-cell Echo DWS measurement scheme, which is the most common implementation of DWS. We present a calibration-free method to blend and merge the Echo and two-cell DWS measurements. Additionally, we demonstrate that using an unbiased fit to the intensity correlation function (ICF) at short times using a regularized inversion algorithm such as CONTIN can substantially improve the quality of the data at very short times. Based on this approach, we further demonstrate that a full correction for the bead and fluid inertia can be applied in a stable manner, further enhancing the microrheology data quality at short times.
@@ II Method
Conventional photon correlation spectroscopy (PCS), as used in DLS or DWS, relies on time-averaging the photon correlation function recorded by a single-photon detector and processed using hardware electronic or software correlator  [16]. Echo-DWS employs a multi-speckle Echo detection scheme to accurately measure slow relaxation dynamics, typically for times equal to or longer than a tenth of a second. Multi-speckle DWS is achieved by illuminating a (e.g. ground-glass) diffuser with a laser and then spinning or oscillating the glass at the desired frequency $1/\tau_{\text{Echo}}$ [42, 44]. The modulated speckle beam is used to illuminate the sample of interest. The resulting, properly ensemble averaged, echo correlation function is combined with a two-cell DWS measurement [36, 41, 17], where the ground glass is displaced slowly to force the decay of the intensity autocorrelation function (ICF) of the sample on typical time scale $\tau_{TC}$. These approaches ensure proper ensemble averaging, achieved by illuminating the sample with a larger number of different speckle fields. Both measurements are then combined to obtain a single ICF spanning correlation times from typically 12.5 ns to ten seconds or more, covering an impressive nine orders of magnitude in relaxation times. The method, as it is currently used, however suffers from two main limitations, both of which we address here.
@@ II.1 Blending and Merging the Two-Cell and Echo Signals
The shortest time accessible to Echo-DWS is determined by the periodicity of the diffuser’s rotation or oscillation. Since this requires mechanically oscillating or rotating an object weighing a few grams with high precision, the minimum accessible echo time is currently limited to just under a tenth of a second. Realizing somewhat shorter times is possible but the procedure tends to be less stable and more prone to drift, or it requires more complex and expensive equipment to provide the required accuracy. In this study, we use an already elevated frequency of 25 Hz, corresponding to the first echo located at $\tau_{\text{Echo}}=0.04$ seconds. Commercial instrumentation, such as the DWS RheoLab (LS Instruments, Switzerland), typically operates with echo-correlation times starting at 0.1 or 0.2 seconds.
To be able to merge the Echo-DWS and the two-cell DWS intensity correlation functions, $g_{2}(t)-1$, the forced decay of the latter must be set to a time similar to or larger than $\tau_{\text{Echo}}$. For comparison, in the original two-cell DWS study by Romer et al. [36], the value of $\tau_{\text{TC}}$ was six times greater than the final data point reported for $g_{2}(t)-1$. Using the multiplication rule proposed in ref. [17], the relation
@@
can be applied to the two-cell data. We note that, for clarity, in this section, we ignore the influence of the experimental coherence factor, $\beta$ and set it to one. By dividing the measured $g_{2}^{\text{TC}}(t)-1$ by $g_{2}^{\text{2nd}}(t)-1$, one can recover the true intensity correlation function $g_{2}(t)-1$, where $g_{2}^{\text{2nd}}(t)-1$ is obtained independently through a long calibration measurement on a solid reference sample, such as a slab made from Teflon or colloidal beads embedded in a solid epoxy matrix.
The maximum time to which the true intensity correlation function can be extracted from TC-DWS reliably depends on the accuracy of the division based on Eq. (1). Unfortunately, the accuracy is often not very high, primarily due to the limited time available for averaging during the recording of $g_{2}^{\text{TC}}(t)-1$. In practice, the forced decay of $g_{2}^{\text{TC}}(t)-1$ exhibits significant variations, and dividing by the reference sample measurement, $g_{2}^{\text{2nd}}(t)-1$, frequently results in visible perturbations of the ICF around $\tau_{\text{Echo}}$. Moreover if two datasets are joined at $\tau_{\text{Echo}}$ without substantial overlap, any deviation of the experimental $g_{2}(t)-1$ from the properly ensemble-averaged value leads to a noticeable kink in the final correlation function used for further analysis, see also Fig. S1. If a greater data overlap is to be achieved using the common method, the measurement time will have to be further extended, which is again not practical.
@@
$\frac{g_{2}^{\text{TC}}(t)-1}{g_{2}^{\text{Echo}}(t)-1}=A\big{(}g_{2}^{\text{2nd}}(t)-1\big{)}=Ae^{-(t/\tau_{\text{TC}})^{2}}.$ (2)
Here, the decay due to the second term is modeled as an exponential function, $e^{-(t/\tau_{\text{TC}})^{2}}$. The $(t/\tau_{\text{TC}})^{2}$-term is attributed to the unidirectional motion of the diffuser.
@@
Using this procedure, we obtain an effective second-cell decay characteristic of the measurement we carried out, removing the need for a calibration sample. We extract $A$ and $\tau_{\text{TC}}$ from a fit to the data over e.g. six data points, and then use $\tau_{\text{TC}}$ to correct $g_{2}^{\text{TC}}(t)-1$ and $A$ to adjust the amplitude of $g_{2}^{\text{Echo}}(t)-1$. Finally, we combine the data over the fitting region, and then append the remaining data points from the echo measurement. This approach ensures a smooth crossover between the two measurements without any discontinuities.
II.2 Analysis of the short-time correlation function decay without bias
[…]
DWS probes $\left[k\left(L/\ell^{\ast}\right)\right]^{2}\left<\Delta\vec{r}^{2}(t)\right>$ over a specific range determined by the measurement precision, where $k=2\pi n/\lambda$ is the wavenumber of scattered light in the host medium with a refractive index $n$. By increasing the optical density $L/\ell^{\ast}$of the sample—either by increasing the cell thickness $L$ or by lowering $\ell^{*}$ through an increase in tracer bead concentration—one can shift the window of accessible mean squared displacements to smaller values (and shorter time scales). However, this comes at the expense of reduced sensitivity to larger displacements.

### v2

@@ Abstract
Diffusing Wave Spectroscopy (DWS) is an extension of standard dynamic light scattering (DLS), applied to soft materials that are turbid or opaque. The propagation of light is modeled using light diffusion, characterized by a light diffusion coefficient that depends on the transport mean free path $\ell^{\ast}$ of the medium. DWS is highly sensitive to small particle displacements or other local fluctuations in the scattering properties and can probe sub-nanometer displacements. Analyzing the motion of beads in a viscoelastic matrix, known as one-bead microrheology, is one of the most common applications of DWS. Despite significant advancements since its invention in the late 1980s, including two-cell and multispeckle DWS, challenges such as merging single- and multispeckle data and limited accuracy for short correlation times persist. Here, we address these issues by improving the two-cell echo DWS scheme. We propose a calibration-free method to blend and merge echo and two-cell DWS data and demonstrate the use of an exponential basis fit to enhance data quality, in particular at very short times. Building on this, we introduce stable corrections for bead and fluid inertia, significantly improving the quality of microrheology data at high frequencies.
@@ I Introduction
Diffusing Wave Spectroscopy (DWS) is a standard light scattering technique used to study the dynamics of colloids, colloid-polymer mixtures, surfactant solutions, and other soft matter systems that display interesting complex relaxation dynamics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Introduced in the late 1980s [14, 15] DWS is an extension of the standard dynamic light scattering (DLS) approach [16]. In contrast to DLS, it is applied to materials that are turbid or opaque. In such materials, light is not scattered only once, as is the case in DLS, but multiple times. The propagation of light is often modeled using light diffusion characterized by a light diffusion coefficient that depends on the transport mean free path $\ell^{\ast}$ of the medium. Coherent light becomes dephased due to the motion of the scatterers in the medium [17, 18, 19], which can be quantified by analyzing the intensity time autocorrelation function (ICF) of the scattered light, analogously to the case of DLS.
@@
DWS is commonly applied to samples such as dense suspensions and emulsions [20, 21, 22, 12, 23, 11, 24]. Alternatively, it can be applied to transparent samples by adding tracer beads that sense the dynamic environment of the medium [25, 26, 7, 5, 27, 28, 29, 22, 13]. The advantage of DWS over DLS is that it can be applied to turbid media that are not accessible to classical DLS, in regimes where multiple scattering suppression techniques such as 3D-cross-correlation fail [30]. Additionally, DWS is more sensitive than DLS to small particle displacements or other local fluctuations in the scattering properties. DWS can probe subnanometer displacements, as demonstrated by Weitz et al. in 1989 shortly after the method was introduced [31]. The high sensitivity of DWS is especially useful when tracer beads are embedded in a viscoelastic medium of interest. DWS can measure viscoelastic solids with elastic moduli of up to 100 kPa [28, 32] and can study relaxations on microsecond timescales, such as Rouse-Zimm and bending modes of polymer dynamics [33, 6, 7].
@@
Analyzing the motion of beads in a viscoelastic matrix is known as (single) bead microrheology [5, 26, 28], which is one of the most common applications of DWS. Pioneering work by Mason, Weitz and others showed that the mean squared displacements of beads $\Delta\vec{r}^{2}(t)>$ can be directly converted into the linear viscoelastic spectrum, or complex modulus $G^{\ast}(\omega)=G^{\prime}(\omega)+iG^{\prime\prime}(\omega)$, of the materials if the radius of the beads is known [25, 34, 35]. DWS provides access to this information from submicrosecond time scales to seconds and minutes, opening opportunities in rheology to study viscoelastic moduli over an enormous range of frequencies with a single measurement [7]. Many reviews of the theory underlying the method, along with numerous examples of its applications, have been published. We refer the reader to the literature for further details [1, 3, 4, 5, 27, 28, 13].
@@
Since the invention of DWS, several substantial improvements have been made to the method. The original approach relied on time averages of the intensity correlation function (ICF), which requires the sample to be in a liquid state. Two-cell DWS, introduced in 2000 [36, 17], and other ensemble averaging approaches [37, 38, 39] provided access to short correlation times for solid nonergodic systems. Multispeckle DWS, using either a camera or the echo scheme, was introduced in the early 2000s and provides access to longer correlation times for highly viscous or viscoelastic solids such as glasses and gels [40, 41, 42, 43, 44, 45, 46, 47, 48]. In 2017, an automated method was developed to determine the mean free paths of optical transport and absorption for absorbing and non-absorbing samples [29]. These advancements have made it possible to access the ICF over a range from typically 10ns to tens of seconds or more, while keeping the measurement time on the order of a few minutes. Only when studying even longer relaxation times does the measurement time need to be further increased.
Despite these significant advances, shortcomings in the implementation of these methods can still lead to errors or reduced data quality. Moreover, access to very short correlation times is hindered by photon shot noise and the need to normalize the ICF, which, in its current implementation, introduces bias.
@@
In the present work, we address these problems and propose two substantial improvements to the implementation of the two-cell echo DWS measurement scheme, which is the most common implementation of DWS. We present a calibration-free method to blend and merge echo and two-cell DWS measurements. Additionally, we demonstrate that using an exponential basis fit directly to the intensity correlation function (ICF) can substantially improve the quality of the data. On the basis of this approach, we further demonstrate that a full correction for the bead and fluid inertia can be applied in a stable manner, further enhancing the microrheology data quality in the important high-frequency regime.
We note that a different approach, diffuse correlation spectroscopy using fast single-photon avalanche diode (SPAD) arrays, has made substantial progress over the past decade. The pixel count of such arrays has increased markedly and, although costs remain high, they have decreased significantly. Sie et al. [49, 50] demonstrated multispeckle diffuse correlation spectroscopy with a $64\times 64$ SPAD array comprising $4096$ pixels and a temporal resolution of $4~\mu\mathrm{s}$, while Wayne et al. reported a $500\times 500$ SPAD camera achieving a full-frame temporal resolution of $10~\mu\mathrm{s}$. Although not yet widely available and still under development, the SPAD approach is becoming increasingly competitive.
@@ II Method
Conventional photon correlation spectroscopy (PCS), as used in DLS or DWS, is based on the time-averaged photon correlation function recorded by a single-photon detector and processed using a hardware electronic or software correlator  [16]. Echo-DWS employs a multispeckle detection scheme to accurately measure the correlation function also in the presence of slow relaxation dynamics, typically for times equal to or longer than tenths of a second. Multispeckle DWS is achieved by illuminating an optical diffuser (e.g., ground glass) with a laser and then spinning or oscillating the glass at the desired frequency $1/\tau_{\text{Echo}}$ [43, 45]. The modulated speckle beam is used to illuminate the sample of interest. We denote by
  • $g_{2}(t)$: the DWS transmission intensity correlation function (ICF) of the sample,
  • $g_{2}^{\mathrm{TC}}(t)$: the ICF measured in a two-cell (TC) configuration, where a slowly moving diffuser (’second cell’) is placed in the beam path together with the sample,
  • $g_{2}^{\mathrm{Echo}}(t)$: the ICF obtained from correlation echoes recorded with a rotating or oscillating diffuser with period $\tau_{\text{Echo}}$,
  • $g_{2}^{\mathrm{2nd}}(t)$: the two-cell ICF for a solid sample, where $g_{2}(t)\equiv 2$, and a slowly moving diffuser. The decay time $\tau_{\text{TC}}$ is controlled by the $\mathrm{2nd}$ cell motion.
In a typical experiment, the properly averaged echo correlation function is combined with a two-cell DWS measurement [36, 42, 17], where the ground glass is slowly displaced to force the decay of the intensity autocorrelation function (ICF) of the sample on a typical time scale $\tau_{TC}$. These approaches ensure proper ensemble averaging, achieved by illuminating the sample with a larger number of different speckle fields. Both measurements are then combined to obtain a single ICF spanning correlation times from typically 12.5 ns to ten seconds or more, covering an impressive nine orders of magnitude in relaxation times. However, the method, as it is currently used, suffers from two main limitations, both of which we address here.
@@ II.1 Blending and Merging the Two-Cell and Echo Signals
The shortest time accessible to Echo-DWS is determined by the periodicity of the rotation or oscillation of the diffuser. Since this requires mechanical oscillation or rotatation of an object weighing a few grams with high precision, the minimum accessible echo time is currently limited to just under a tenth of a second. Realizing somewhat shorter times is possible, but the procedure tends to be less stable and more prone to drift, requires more startup time for stabilization, or requires more complex and expensive equipment to provide the required accuracy. In this study, we used an already elevated frequency of 25 Hz, corresponding to the first echo located at $\tau_{\text{Echo}}=0.04$ seconds. Commercial instrumentation, such as the DWS RheoLab (LS Instruments, Switzerland), typically operates with echo-correlation times that start at $0.1$- seconds.
To be able to merge the Echo-DWS and the two-cell DWS intensity correlation functions, $g_{2}(t)-1$, the forced decay of the latter must be set to a time similar to or greater than $\tau_{\text{Echo}}$. For comparison, in the original two-cell DWS study by Romer et al. [36], the value of $\tau_{\text{TC}}$ was six times greater than the final data point reported for $g_{2}(t)-1$. Using the multiplication rule proposed in ref. [17], the relation
@@
can be applied to the two-cell data. By dividing the measured $g_{2}^{\text{TC}}(t)-1$ by $g_{2}^{\text{2nd}}(t)-1$, we recover the true intensity correlation function $g_{2}(t)-1$, where $g_{2}^{\text{2nd}}(t)-1$ is obtained independently through a long calibration measurement on a solid reference sample, such as a slab made of Teflon or colloidal beads embedded in a solid epoxy matrix. We note that for clarity, in this section we ignore the influence of the experimental coherence factor $\beta$ and set it to one.
The maximum time to which the true intensity correlation function can be reliably extracted from TC-DWS depends on the precision of the division based on Eq. (1). Unfortunately, the accuracy is often not very high, primarily due to the limited time available to average during the recording of $g_{2}^{\text{TC}}(t)-1$. For example, for $\tau_{\text{TC}}=0.3$, with a total measurement time of 300 seconds this implies that the time average of $g_{2}^{\mathrm{TC}}(t)-1$ covers $\sim 300/0.3=1000$ statistically independent speckle realizations. The statistical noise scales as $1/\sqrt{1000}\simeq 3\%$ [48]; see also Table S1. Thus, in practice, the forced decay of $g_{2}^{\text{TC}}(t)-1$ exhibits significant variations, and dividing by the measurement of the reference sample, $g_{2}^{\text{2nd}}(t)-1$, frequently results in visible perturbations of the ICF around $\tau_{\text{Echo}}$. Moreover, if $g_{2}(t)-1$ is joined with $g^{\text{Echo}}_{2}(t)-1$ at $\tau_{\text{Echo}}$ without data overlap, any deviation of the experimental $g_{2}(t)-1$ from the properly ensemble-averaged value leads to a noticeable kink in the final correlation function used for further analysis; see also Fig. S1. If data overlap is to be achieved using the standard approach, one would need to increase $\tau_{\mathrm{TC}}$, requiring a manifold extension of the measurement time, which is not practical.
@@
$\frac{g_{2}^{\text{TC}}(t)-1}{g_{2}^{\text{Echo}}(t)-1}=A\big(g_{2}^{\text{2nd}}(t)-1\big)=Ae^{-(t/\tau_{\text{TC}})^{p}}.$ (2)
The decay caused by the second cell, $g_{2}^{\text{2nd}}(t)-1$ , is modeled as an exponential function, $e^{-(t/\tau_{\text{TC}})^{p}}$. Due to the unidirectional motion of the diffuser $p=2$ and $\tau_{\text{TC}}=\text{const.}$, see Fig. S4. However, for finite measurement times, both values, $A$ and $p$, can fluctuate around the mean value. For the examples shown in the main text that study a viscoelastic liquid, we find that $p=2=\text{const.}$ can be used, while for the viscoelastic solid reported in the supplemental material, fluctuations in $p$ become noticeable, and therefore both $\tau_{\text{TC}}$ and $p$ have to be treated as adjustable parameters.
@@
Using this procedure, we obtain an effective second-cell decay characteristic of the measurement we carried out, removing the need for a calibration sample. We extract $A$ and $\tau_{\text{TC}}$ from a fit to the data over, e.g., six data points, and then use $\tau_{\text{TC}}$ to correct $g_{2}^{\text{TC}}(t)-1$ and $A$ to adjust the amplitude of $g_{2}^{\text{Echo}}(t)-1$. Finally, we combine the data over the fitting region and then append the remaining data points from the echo measurement. This approach ensures a smooth crossover between the two measurements without discontinuities.
II.2 Analysis of the short-time correlation function decay
One of the key features of DWS is the ability to measure tiny displacements of particles or other local rearrangements. DWS enables measurements of sub-diffusive motion on nanometer length scales using visible light as a probe. In DWS-based microrheology, relaxations on (sub)microsecond time scales correspond to frequencies in the viscoelastic spectra of $10^{6}\,\text{rad/s}$ and above, which no other laboratory technique can measure. For comparison, piezo-driven squeeze-flow devices can reach frequencies up to approximately $10^{4}\,\text{rad/s}$ [51]. Torsional resonance oscillators extend the accessible frequency range into the $10^{5}\,\text{rad/s}$ regime [52]; however, rheological measurements at such high frequencies are often limited by experimental constraints and are therefore impractical in most routine applications. Displacements of particles on these time scales are sub-Brownian, which means they are increasingly affected by the finite inertia of the tracer particles or beads and the surrounding fluid, which we will discuss in the next section. First, we focus on the accurate measurement of the short-time decay. DWS provides $\left[k\left(L/\ell^{\ast}\right)\right]^{2}\left<\Delta\vec{r}^{2}(t)\right$> over a specific range determined by the precision of the measurement, where $k=2\pi n/\lambda$ is the wavenumber of scattered light in the host medium with a refractive index $n$. By increasing the optical density $L/\ell^{\ast}$ of the sample - either by increasing the cell thickness $L$ or by lowering $\ell^{*}$ through an increase in the concentration of tracer beads - one can shift the window of accessible mean squared displacements to smaller values (and shorter time scales). However, this comes at the expense of reduced sensitivity to larger displacements.